OpenStudy (khally92):

need help. find the first derivative (1-2x)^2(x-3). How am i to go about it, do i need to expand the first bracket.

OpenStudy (anonymous):

you can expand the first bracket if you want. I would not expand and use the chain rule on it instead.

OpenStudy (anonymous):

i would expand the whole thing \[(1-2x)^2(x-3)=4x^2-16x^2+13x-3\] and then it is real easy

OpenStudy (khally92):

i used the chain rule as well, could u please help verify the answer, this is what i got -4(1-2x)(x-3)+(1-2x)^2. i will xpand it and c if i get d same answer.

OpenStudy (anonymous):

I was going to say expanding is fine too...ill check

OpenStudy (anonymous):

if you do it that way you have to use the product rule too

OpenStudy (anonymous):

should all work out the same since all the short cut rules are consistent but for my money if you are going to have to expand anyway, why not to it first?

OpenStudy (anonymous):

@satellite73 agreed

OpenStudy (anonymous):

Yeah I get the same

OpenStudy (khally92):

no way.... u cannot get the same when u expand the whole thing. i expand n got 13-16x-12x^2-3

OpenStudy (khally92):

well if u got the same as well then how can i go about findind the second derivative.

OpenStudy (anonymous):

better to expand the original function first then differentiate twice

OpenStudy (anonymous):

easier to deal with

OpenStudy (khally92):

so please verify my answer when i expanded the whole equation 13-16x-12x^2-3

OpenStudy (anonymous):

(1-2x)^2(x-3)=\[(4x^2-4x+1))(x-3)=4x^3-12x^2-4x^2+12x+x-3\]\[4x^3-16x^2+13x-3\]

OpenStudy (anonymous):

now differentiate

OpenStudy (anonymous):

\[f'(x)=12x^2-32x+13\]

OpenStudy (anonymous):

\[f''(x)=24x-32\]

OpenStudy (khally92):

-4x^3+8x^2+13x-3

OpenStudy (khally92):

thats my ans. for first derivative.

OpenStudy (anonymous):

where do you get -4x^3

OpenStudy (khally92):

if u expand (1-2x)^2 you will get 1-4x-4x^2 ryt?

OpenStudy (anonymous):

wrong. 4x^2-4x+1

OpenStudy (khally92):

let me go bck again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(-2x)^2=4x^2

OpenStudy (khally92):

well well well i got 12x^2-32x+13. as the 1st derivative

OpenStudy (khally92):

is dat correct?

OpenStudy (khally92):

hello eseidi are u ther?

OpenStudy (anonymous):

yep, that should be right.